OpenStudy (anonymous):

If u and v are the vectors below, find the vector w whose tail is at the point halfway from the tip of v to the tip of u and whose head is at the point halfway from the tip of u to the tip of u-v. u = (-2,-4,-5) v = (-3, 5, -1) answer: (2, -7, -3/2)

OpenStudy (anonymous):

tail = 1/2(u-v) = 1/2(1,-9,-4) = (1/2,-9/2,-2) head = 1/2(u-u+v) = 1/2(-3,5,-1) = (-3/2, 5/2,-1/2) w = (-2,7,1/2)

OpenStudy (anonymous):

oh, wrong reading, head = 1/2(u-v-u) = 1/2(3,-5,1) = (3/2,-5/2,1/2) so, w = head - tail = (1,2,5/2)